# HOLESOM: Constraining the Properties of Slowly-Accreting Massive Black Holes with Self-Organizing Maps

### v3

@@ Abstract
Accreting massive black holes (MBHs, with $M_{\bullet}>10^{3}\,{\rm M_{\odot}}$) are known for their panchromatic emission, spanning from radio to gamma rays. While MBHs accreting at significant fractions of their Eddington rate are readily detectable, those accreting at much lower rates in radiatively inefficient modes often go unnoticed, blending in with other astrophysical sources. This challenge is particularly relevant for gas-starved MBHs in external galaxies and those possibly wandering in the Milky Way. We present HOLESOM, a machine learning-powered tool based on the Self-Organizing Maps (SOMs) algorithm, specifically designed to identify slowly-accreting MBHs using sparse photometric data. Trained on a comprehensive set of $\sim 20,000$ spectral energy distributions (SEDs), HOLESOM can (i) determine if the photometry of a source is consistent with slowly-accreting MBHs and (ii) estimate its black hole mass and Eddington ratio, including uncertainties. We validate HOLESOM through extensive tests on synthetic data and real-world cases, including Sagittarius A* (Sgr A*), demonstrating its effectiveness in identifying slowly-accreting MBHs. Additionally, we derive analytical relations between radio and X-ray luminosities to further constrain physical parameters. The primary strength of HOLESOM lies in its ability to accurately identify MBH candidates, which can then be targeted for follow-up photometric and spectroscopic observations. Fast and scalable, HOLESOM offers a robust framework for automatically scanning large multi-wavelength datasets, making it a valuable tool for unveiling hidden MBH populations in the local Universe.
@@ 1 Introduction
where $\dot{M}_{\rm Edd}$ is the mass accretion rate at which a relativistic thin accretion disk (Novikov & Thorne, 1973) would radiate at the Eddington luminosity (see, e.g., the discussion in Pacucci & Narayan 2024), and $\dot{M}$ is the mass accretion rate onto the MBH. In general, the radiated luminosity $L$ is related to the accretion rate $\dot{M}$ via $L=\epsilon\dot{M}c^{2}$, where $\epsilon$ is the matter-to-energy radiative efficiency. In the standard, thin-disk accretion scenario (Shakura & Sunyaev, 1976), valid for $10^{-2}\lesssim\,{f_{\rm Edd}}\lesssim 1$, the matter-to-energy radiative efficiency is $\epsilon\approx 10\%$ and depends on the black hole spin (Bardeen, 1970; Thorne, 1974). Highly-accreting MBHs are also observed at high redshifts, with the farthest confirmed AGN now standing at $z=10.6$ (Maiolino et al., 2023), although the sample completeness is much lower in these cases.
On the contrary, even in the local Universe, the population of MBHs accreting at very low rates ($\,{f_{\rm Edd}}<10^{-2}$) is sparsely studied. For such low rates, the radiative efficiency $\epsilon$ decreases dramatically, possibly by orders of magnitude, as we enter the domain of advection-dominated accretion flows (ADAF, Narayan & Yi 1994; Abramowicz et al. 1995; Narayan & McClintock 2013). As the conditions for large accretion rates are rare, especially at low redshift (Power et al., 2010), most MBHs likely accrete in ADAF mode (Narayan et al., 2022). An example of such objects is the SMBH at the center of the Milky Way (MW): Sagittarius A* (Sgr A*, in short). Recently imaged by the Event Horizon Telescope collaboration (Event Horizon Telescope Collaboration et al., 2022), this SMBH is accreting at $\,{f_{\rm Edd}}\sim 10^{-7}$ (see, e.g., Ressler et al. 2018) and was initially detected by studying the dynamics of the stars in its immediate vicinity (Ghez et al., 2000; Genzel et al., 2000).
@@
For these reasons, slowly-accreting MBHs in local galaxies can be selected only via a massively multi-wavelength approach. Such objects can be discerned from other contaminant categories only by collecting numerous photometric points in different bands and comparing this photometry with ADAF SEDs generated either analytically (see, e.g., Mahadevan 1997; Pesce et al. 2021), or with general-relativistic magneto-hydrodynamic (GRMHD) simulations of the accretion flow, down to a few gravitational radii (see, e.g., Dolence et al. 2009). The use of advanced computational techniques can lead to a better identification and characterization of such sources.
Machine learning (ML) techniques offer an exciting prospect for selecting and characterizing the physical properties of slowly accreting MBHs in the local Universe. Once a ML technique is trained with the most extensive available set of ADAF SEDs, it learns the patterns between observables and physical underlying parameters, such that it can be deployed to analyze multi-wavelength data fields pertaining to the local Universe. Such codes are easily scalable and can be fed massive amounts of data once proper source-matching techniques between catalogs are used (see, e.g., the software LYRA, Peca et al. 2021).
In this study, we present HOLESOM, a software tool based on the unsupervised ML technique of Self-Organizing Maps. This software represents the first attempt at “black hole cartography”, aimed at identifying and mapping slowly-accreting MBHs in the correct location of the $M_{\bullet}$ vs. $\,{f_{\rm Edd}}$ parameter space. This study is organized as follows. In Sec. 2, we present the analytical SEDs used to train our model. Then, Sec. 3 describes the ML methods we employ. In Sec. 4, we present our results. We test our software with artificial (Sec.4.1) and real-world examples (Sec.4.2). Then, we show that the photometry of slowly-accreting MBHs can be effectively disentangled from that of stellar-mass compact objects (Sec.4.3). We also present relationships between radio luminosity, X-ray luminosity, black hole mass and Eddigton rate (Sec.5). We conclude with Sec. 6 discussing future applications of HOLESOM.
@@ 3.1 Self-Organizing Maps
We develop HOLESOM on the basis of the ML algorithm named Self-Organizing Maps (SOMs) (Kohonen, 1982). SOMs are a type of unsupervised artificial neural network (see, e.g., Baron 2019 for a review) that reduces a high-dimensional parameter space to a lower-dimensional representation while preserving the original topology. In this way, similar objects in the high-dimensional parameter space will be grouped in the final map, maintaining the intrinsic structure of the input data. In this study, we reduce the dimensionality to $2$, as this choice facilitates visualization. We use the publicly-available code SomPY [^1] https://github.com/sevamoo/SOMPY library (Moosavi et al., 2014a) to construct and train our SOM.
Each pixel of the 2D map is characterized by a weight vector corresponding to a position in the higher-dimensional parameter space of the input data, normalized to the unit variance with a mean of zero. The weight vectors are initialized through principal component analysis (PCA; Chatfield & Collins 1980) and then updated iteratively to minimize the Euclidean distance to the input data and reduce the quantization error.
@@ 3.1.1 Training and Visualization
These input features were carefully selected among all the values of the luminosities $\nu L_{\nu}$ and all the possible SED slopes (see Section 2) to ensure homogeneous frequency coverage from radio to X-rays and strong correlation with $M_{\bullet}$ and $\,{f_{\rm Edd}}$. Specifically, we computed the distance correlation between $M_{\bullet}$ and $\,{f_{\rm Edd}}$ with the luminosities as well as the SED slopes across all possible combinations (i.e., $1326$) of two frequencies.
For each pixel, we compute the median values of $\,{f_{\rm Edd}}$ and $M_{\bullet}$. The top panel of Fig. 2 shows the fully trained SOM color-coded by median $\log_{10}\,{f_{\rm Edd}}$, with values smoothly increasing vertically from top to bottom. In contrast, median $\log_{10}M_{\bullet}$ values progressively increase leftwards, as shown by the shades of gray contours with values indicated in the color bar on the left.
@@ 4 Results
In this Section, we apply the ML method described thus far, adopting a progressive approach. First, we test how well we can constrain the main physical parameters of a putative source if data is available only in a very limited set of frequencies (Sec. 4.1). This, of course, will be the case with any real-world application. Second, we test HOLESOM with the archetypal example of an ADAF MBH: Sgr A (Sec. 4.2). In doing this, we highlight why this source is peculiar and possibly not representative of the majority of sources to which our methodology can be applied. Then, in Sec. 4.3 we show that the MBHs accreting in ADAF mode can be effectively disentangled from stellar-mass accretors (e.g., stellar-mass black holes, neutron stars, pulsars).
@@ 4.1 Evaluating $\,{f_{\rm Edd}}$ and $M_{\bullet}$ ranges with limited data
We now test the fully-trained SOM to investigate how it performs in real-world scenarios. In an ideal situation, a given source possesses a large wealth of multi-wavelength data, spanning from radio to X-ray. In most cases, however, only a limited amount of data can be securely associated with a given source, with a number of data points that is typically lower than the $10$ input features with which our code was trained. The relevant question is: how well the physical parameters of the MBH can be constrained if one or more multi-wavelength data points are missing?
@@
Hence, our goal is to investigate whether, with scarce data we can: (i) correctly classify the mass category of a MBH and (ii) constrain its Eddington ratio within the upper end of the ADAF range (i.e., $\,{f_{\rm Edd}}\sim 10^{-2}-10^{-3}$), or at severely low rates (i.e., $\,{f_{\rm Edd}}\ll 10^{-3}$). We first discuss the ideal scenario with no missing observations. Next, we explain how to estimate parameters when data is missing, beginning with the case of a single missing photometric point and concluding with scenarios of multiple missing observations.
@@
For a more realistic scenario, we remove one photometric band at a time. We then recover that specific missing information following the method detailed in La Torre et al. (2024), to which the interested reader is referred to. In summary, this method consists of randomly drawing $5000$ new values of the missing feature from the distribution of the rest of the training sample, under the assumption that the missing value falls within the same distribution as the sample and it is not a peculiar case. Then, given the new set of input features, we project the MBH onto the SOM, creating a 2D likelihood surface of the possible pixel occupied by the MBH. From that, we extract the probability distribution function of $\,{f_{\rm Edd}}$ and $M_{\bullet}$ for that MBH, given the recovered missing data.
@@
For $\,{f_{\rm Edd}}$, we observe that the ranges are consistent with the true value, especially in the upper sub-Eddington regime around 10-3-10-2 (i.e., Figure 9 in the Appendix), where $\,{f_{\rm Edd}}$ is mainly affected by IR, optical, UV, and X-ray observations. This result also holds in the low end of $\,{f_{\rm Edd}}$ around $10^{-8}-10^{-7}\,{\rm M_{\odot}}yr^{-1}$ (i.e., the top two panels in Figure 8 in the Appendix). Even with missing data, we can constrain $\,{f_{\rm Edd}}$ to be in the upper or lower ranges in the sub-Eddington regime. This result slightly differs for a MBH in an intermediate sub-Eddington regime (i.e., $\sim 10^{-5}$), as shown in the bottom row of Figure 8, for a MBH in pixel 3. Here, radio observations start to affect the process of constraining $\,{f_{\rm Edd}}$; specifically, recovering the missing luminosity at $950$ GHz would place this object in a deeper sub-Eddington regime.
Recovering missing observations yields consistent ranges of $M_{\bullet}$ with the true values for cases at the high end of the black hole mass (i.e., $\sim 10^{7}\,{\rm M_{\odot}}$). Unlike $\,{f_{\rm Edd}}$, radio and MIR luminosities mainly influence the mass range. This pattern also holds in the opposite regime, i.e., at the low mass end (pixel 2 and 5 in Figure 8), where recovering missing observations still yields consistent mass ranges compared to the true value. However, missing luminosity at $5$ GHz and/or at $8.6\,\mu$m could result in mass regimes differing by $\sim 1-2$ orders of magnitude. For the intermediate mass case (pixel 3), the mass ranges are consistent with the true value within the margin of error.
@@
Figure 3: For a test MBH located in pixel 4, we illustrate the values of Log $\,{f_{\rm Edd}}$ (left) and $M_{\bullet}$ (right) estimated with HOLESOM, in various cases of data availability. The numbers in parenthesis indicate how many of the 10 bands are available. In the top rows, labeled ‘All’, we show the true values of the test MBH, marked by a black cross. The black square indicates, when all bands are available, the median of the parameter distribution within pixel 4, with the grey line showing the standard deviation of that distribution. In the lower rows, labeled ‘Missing’, we show the parameter ranges after recovering one or multiple missing observations (Section 4.1). In these cases, the red circle marks the peak of the probability distribution function for the estimated parameter after recovering missing data, while the grey line again represents the standard deviation. Additional test MBHs from other pixels are shown in the Appendix A.
In conclusion, in the ideal scenario of no missing data (i.e., $\sim 10$ photometric data points), $\,{f_{\rm Edd}}$ and $M_{\bullet}$ are estimated with high accuracy. In case of missing X-ray data or, more generally, a single-band observation, HOLESOM remains effective in correctly classifying MHBs within reasonable mass and accreting ADAF ranges. In the case of numerous missing data points, the uncertainty is higher, but HOLESOM still remains effective in correctly identifying a slowly-accreting MBH and approximating its mass and Eddington ratio.

### v4

@@ Abstract
Accreting massive black holes (MBHs, with $M_{\bullet}>10^{3}\,{\rm M_{\odot}}$) are known for their panchromatic emission, spanning from radio to gamma rays. While MBHs accreting at significant fractions of their Eddington rate are readily detectable, those accreting at much lower rates in radiatively inefficient modes often go unnoticed, blending in with other astrophysical sources. This challenge is particularly relevant for gas-starved MBHs in external galaxies and those possibly wandering in the Milky Way. We present HOLESOM [^1] HOLESOM is publicly available at:
https://github.com/valentinalatorre/holesom
, a machine learning-powered tool based on the Self-Organizing Maps (SOMs) algorithm, specifically designed to identify slowly-accreting MBHs using sparse photometric data. Trained on a comprehensive set of $\sim 20,000$ spectral energy distributions (SEDs), HOLESOM can (i) determine if the photometry of a source is consistent with slowly-accreting MBHs and (ii) estimate its black hole mass and Eddington ratio, including uncertainties. We validate HOLESOM through extensive tests on synthetic data and real-world cases, including Sagittarius A (Sgr A), demonstrating its effectiveness in identifying slowly-accreting MBHs. Additionally, we derive analytical relations between radio and X-ray luminosities to further constrain physical parameters. The primary strength of HOLESOM lies in its ability to accurately identify MBH candidates, which can then be targeted for follow-up photometric and spectroscopic observations. Fast and scalable, HOLESOM offers a robust framework for automatically scanning large multi-wavelength datasets, making it a valuable tool for unveiling hidden MBH populations in the local Universe.
@@ 1 Introduction
where $\dot{M}_{\rm Edd}$ is the mass accretion rate at which a relativistic thin accretion disk (Novikov & Thorne, 1973) would radiate at the Eddington luminosity (see, e.g., the discussion in Pacucci & Narayan 2024), and $\dot{M}$ is the mass accretion rate onto the MBH. In general, the radiated luminosity $L$ is related to the accretion rate $\dot{M}$ via $L=\epsilon\dot{M}c^{2}$, where $\epsilon$ is the matter-to-energy radiative efficiency. In the standard, thin-disk accretion scenario (Shakura & Sunyaev, 1976), valid for $10^{-2}\lesssim\,{f_{\rm Edd}}\lesssim 1$, the matter-to-energy radiative efficiency is $\epsilon\approx 10\%$ and depends on the black hole spin (Bardeen, 1970; Thorne, 1974). Highly-accreting MBHs are also observed at high redshifts, with the farthest confirmed AGN now standing at $z=10.6$ (Maiolino et al., 2024), although the sample completeness is much lower in these cases.
On the contrary, even in the local Universe, the population of MBHs accreting at very low rates ($\,{f_{\rm Edd}}<10^{-2}$) is sparsely studied. For such low rates, the radiative efficiency $\epsilon$ decreases dramatically, possibly by orders of magnitude, as we enter the domain of advection-dominated accretion flows (ADAF, Narayan & Yi 1994; Abramowicz et al. 1995; Narayan & McClintock 2013). As the conditions for large accretion rates are rare, especially at low redshift (Power et al., 2010), most MBHs likely accrete in ADAF mode (Narayan et al., 2022). An example of such objects is the SMBH at the center of the Milky Way (MW): Sagittarius A (Sgr A, in short). Recently imaged by the Event Horizon Telescope collaboration (Event Horizon Telescope Collaboration et al., 2022), this SMBH is accreting at $\,{f_{\rm Edd}}\sim 10^{-7}$ (see, e.g., Ressler et al. 2018) and was initially detected by studying the dynamics of the stars in its immediate vicinity (Ghez et al., 2000; Genzel et al., 2000).
@@
For these reasons, slowly accreting MBHs in local galaxies can be selected only via a massively multi-wavelength approach. Such objects can be discerned from other contaminant categories only by collecting numerous photometric points in different bands and comparing this photometry with ADAF SEDs generated either analytically (see, e.g., Mahadevan 1997; Pesce et al. 2021), or with general-relativistic magneto-hydrodynamic (GRMHD) simulations of the accretion flow, down to a few gravitational radii (see, e.g., Dolence et al. 2009). Advanced computational techniques can lead to better identification and characterization of such sources.
Machine learning (ML) techniques offer an exciting prospect for selecting and characterizing the physical properties of slowly accreting MBHs in the local Universe. Once an ML technique is trained with the most extensive available set of ADAF SEDs, it learns the patterns between observables and physical underlying parameters such that it can be deployed to analyze multi-wavelength data fields pertaining to the local Universe. Such codes are easily scalable and can be fed massive amounts of data once proper source-matching techniques between catalogs are used (see, e.g., the software LYRA, Peca et al. 2021).
In this study, we present HOLESOM, a software tool based on the unsupervised ML technique of Self-Organizing Maps. This software represents the first attempt at “black hole cartography”, aimed at identifying and mapping slowly-accreting MBHs in the correct location of the $M_{\bullet}$ vs. $\,{f_{\rm Edd}}$ parameter space. This study is organized as follows. In Sec. 2, we present the analytical SEDs used to train our model. Then, Sec. 3 describes the ML methods we employ. In Sec. 4, we present our results. We test our software with artificial (Sec. 4.1) and real-world examples (Sec. 4.2). Then, we show that HOLESOM can automatically identify the photometry of slowly-accreting MBHs candidates while rejecting different astrophysical sources (Sec. 4.3). Sec. 5 gives as overview of the procedure for utilizing HOLESOM. We also present relationships between radio luminosity, X-ray luminosity, black hole mass, and Eddington rate (Sec. 6). We conclude with Sec. 7 discussing future applications of HOLESOM.
@@ 3.1 Self-Organizing Maps
We develop HOLESOM on the basis of the ML algorithm named Self-Organizing Maps (SOMs) (Kohonen, 1982), and make it publicly available at https://github.com/valentinalatorre/holesom.
SOMs are a type of unsupervised artificial neural network (see, e.g., Baron 2019 for a review) that reduces a high-dimensional parameter space to a lower-dimensional representation while preserving the original topology. In this way, similar objects in the high-dimensional parameter space will be grouped in the final map, maintaining the intrinsic structure of the input data. In this study, we reduce the dimensionality to $2$, as this choice facilitates visualization. We use the publicly-available code SomPY [^2] https://github.com/sevamoo/SOMPY library (Moosavi et al., 2014) to construct and train our SOM.
Each pixel of the 2D map is characterized by a weight vector corresponding to a position in the higher-dimensional parameter space of the input data, normalized to the unit variance with a mean of zero. The weight vectors are initialized through principal component analysis (PCA; Chatfield & Collins 1980) and then updated iteratively to minimize the Euclidean distance to the input data and reduce the quantization error. For each input data point, the cell whose weight vector has the smallest Euclidean distance to it is identified as the Best Match Unit (BMU).
We adopt an unsupervised approach with SOM rather than a supervised method for several reasons. First, supervised models require extensive hyperparameter tuning, model parameter and optimization choices that significantly impact model performance and computational time. In contrast, SOM operates directly in the data space with a simpler set of hyperparameters and does not require strong prior assumptions about model complexity. Second, we desire uncertainty in our results rather than a point estimate. Obtaining uncertainty estimates with supervised methods typically requires assuming an underlying data distribution (e.g. Gaussian, multimodal). SOM is non-parametric and can be used to construct a posterior probability distribution function (PDF) entirely from the data itself (López-Rubio, 2010; Hemmati et al., 2019; La Torre et al., 2024). Third, SOM provides an intuitive 2D representation of the data, enabling the identification of relationships and clusters that might not be apparent in a supervised framework. Finally, with SOM, we can easily implement methods for handling missing data without retraining the model (La Torre et al., 2024), as discussed in Section 4.1. This is crucial when dealing with sparse photometric observations, which can vary across test sources, as is expected for MBH data. In summary, SOM can provide visualization (Sec. 3.1.1), parameter estimation with PDFs (Sec. 4.1, 4.2), and classification (Sec. 4.3) in a single framework, offering a more interpretable and flexible model with lower computational complexity.
@@ 3.1.1 Training and Visualization
These input features were carefully selected based on two main criteria. First, we aimed to balance homogeneous spectral coverage with practical constraints of ML applications and observational limitations. While a continuous sampling approach is theoretically ideal, it is impractical for MBH applications, where continuous spectral coverage is rarely available. Indeed, in reality, MBHs have only a few detections, and adding more features would only exacerbate the issue of missing data, as ML models require consistency between training and test sets. Therefore, we select a minimal yet representative set of features that ensures homogeneous coverage across the SED while balancing the constraints of ML algorithms and the observational limitations of MBHs. By adopting a representative and widely spaced sampling of the SED, we mitigate the impact of missing data–an expected issue with real data, though we do not know a priori which bands will be missing for a given source–while preserving meaningful physical information. The treatment of missing data is further discussed in Section 4.1.
@@ 50 GHz $5\times 10^{10}$
Second, we prioritize features most sensitive to physical parameters such as $M_{\bullet}$ and $\,{f_{\rm Edd}}$. To identify these, we computed the distance correlation between $M_{\bullet}$ and $\,{f_{\rm Edd}}$ with all the luminosities $\nu L_{\nu}$ across 54 frequencies from radio to X-ray, as well as all possible two-frequency combinations of SED slopes (1,326 combinations). After selecting the features most strongly correlated with these parameters, we trained multiple SOMs with different feature sets. The final 10 features reported in Table 1 led to the smallest dispersions in $M_{\bullet}$ and $\,{f_{\rm Edd}}$ within SOM cells.
We compute the median values of $\,{f_{\rm Edd}}$ and $M_{\bullet}$ for each pixel. The top panel of Fig. 2 shows the fully trained SOM color-coded by median $\log_{10}\,{f_{\rm Edd}}$, with values smoothly increasing vertically from top to bottom. In contrast, median $\log_{10}M_{\bullet}$ values progressively increase leftwards, as shown by the shades of gray contours with values indicated in the color bar on the left.
@@ 4 Results
In this Section, we apply the ML method described thus far, adopting a progressive approach. First, we test how well we can constrain the main physical parameters of a putative source if data is available only in a very limited set of frequencies (Sec. 4.1). This, of course, will be the case with any real-world application. Second, we test HOLESOM with the archetypal example of an ADAF MBH: Sgr A (Sec. 4.2). In doing this, we highlight why this source is peculiar and possibly not representative of the majority of sources to which our methodology can be applied. Then, in Sec. 4.3, we show that HOLESOM can effectively identify MBHs accreting in ADAF mode while filtering out sources with non-MBH photometry.
@@ 4.1 Evaluating $\,{f_{\rm Edd}}$ and $M_{\bullet}$ ranges with limited data
We now test the fully-trained SOM to investigate its performance in real-world scenarios. In an ideal situation, a given source possesses a large wealth of multi-wavelength data, spanning from radio to X-ray. In most cases, however, only a limited amount of data can be securely associated with a given source, with several data points typically lower than the $10$ input features with which our code was trained. The relevant question is: How well can the physical parameters of the MBH be constrained if one or more multi-wavelength data points are missing?
@@
Hence, our goal is to investigate whether, with scarce data, we can: (i) correctly classify the mass category of a MBH and (ii) constrain its Eddington ratio within the upper end of the ADAF range (i.e., $\,{f_{\rm Edd}}\sim 10^{-2}-10^{-3}$), or at severely low rates (i.e., $\,{f_{\rm Edd}}\ll 10^{-3}$). We first discuss the ideal scenario with no missing observations. Next, we explain how to estimate parameters when data is missing, beginning with the case of a single missing photometric point and concluding with scenarios of multiple missing observations.
@@
For a more realistic scenario, we remove one photometric band at a time. We then recover that specific missing information following the method detailed in La Torre et al. (2024), to which the interested reader is referred. In summary, this method consists of randomly drawing $5000$ new values of the missing feature from the distribution of the rest of the training sample. This approach is justified as we can reasonably assume that the missing value, due to non-detection, falls within the same distribution as the training sample, which contains values for the MBH population, and it is not an outlier or a peculiar case. Then, given the new set of input features, we project the MBH onto the SOM, creating a 2D likelihood surface of the possible pixel occupied by the MBH. From that, we extract the probability distribution function of $\,{f_{\rm Edd}}$ and $M_{\bullet}$ for that MBH, given the recovered missing data.
@@
Figure 3: For a test MBH located in pixel 4, we illustrate the values of Log $\,{f_{\rm Edd}}$ (left) and $M_{\bullet}$ (right) estimated with HOLESOM, in various cases of data availability. The numbers in parentheses indicate how many of the 10 bands are available. In the top rows, labeled ‘All’, we show the true values of the test MBH, marked by a black cross. The black square indicates, when all bands are available, the median of the parameter distribution within pixel 4, with the grey line showing the standard deviation of that distribution. In the lower rows, labeled ‘Missing’, we show the parameter ranges after recovering one or multiple missing observations (Section 4.1). In these cases, the red circle marks the peak of the probability distribution function for the estimated parameter after recovering missing data, while the grey line again represents the standard deviation. Additional test MBHs from other pixels are shown in the Appendix A.
For $\,{f_{\rm Edd}}$, we observe that the ranges are consistent with the true value, especially in the upper sub-Eddington regime around 10-3-10-2 (i.e., Figure 10 in the Appendix), where $\,{f_{\rm Edd}}$ is mainly affected by IR, optical, UV, and X-ray observations. This result also holds in the low end of $\,{f_{\rm Edd}}$ around $10^{-8}-10^{-7}\,{\rm M_{\odot}}yr^{-1}$ (i.e., the top two panels in Figure 9 in the Appendix). Even with missing data, we can constrain $\,{f_{\rm Edd}}$ to be in the upper or lower ranges in the sub-Eddington regime. This result slightly differs for a MBH in an intermediate sub-Eddington regime (i.e., $\sim 10^{-5}$), as shown in the bottom row of Figure 9, for a MBH in pixel 3. Here, radio observations start to constrain $\,{f_{\rm Edd}}$; specifically, recovering the missing luminosity at $950$ GHz would place this object in a deeper sub-Eddington regime.
Recovering missing observations yields consistent ranges of $M_{\bullet}$ with the true values for cases at the high end of the black hole mass (i.e., $\sim 10^{7}\,{\rm M_{\odot}}$). Unlike $\,{f_{\rm Edd}}$, radio and MIR luminosities mainly influence the mass range. This pattern also holds in the opposite regime, i.e., at the low mass end (pixel 2 and 5 in Figure 9), where recovering missing observations still yields consistent mass ranges compared to the true value. However, missing luminosity at $5$ GHz and/or $8.6\,\mu$m could result in mass regimes differing by $\sim 1-2$ orders of magnitude. For the intermediate mass case (pixel 3), the mass ranges are consistent with the true value within the margin of error.
@@
In conclusion, in the ideal scenario of no missing data (i.e., $\sim 10$ photometric data points), $\,{f_{\rm Edd}}$ and $M_{\bullet}$ are estimated with high accuracy. In case of missing X-ray data or, more generally, a single-band observation, HOLESOM remains effective in correctly classifying MHBs within reasonable mass and accreting ADAF ranges. The uncertainty is higher in the case of numerous missing data points, but HOLESOM remains effective in correctly identifying a slowly-accreting MBH and approximating its mass and Eddington ratio.